40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3933     Entry time: Tue Nov 16 15:32:18 2010
Author: valera 
Type: Update 
Category: Electronics 
Subject: OSEM noise at the output of the satellite box 

 I measured the SRM OSEM (no magnets at the moment) noise out of the satellite box with a SRS785 spectrum analyzer. I inserted a break out board into the cable going from the satellite box to the whitening board. The transimpedances of the SRM OSEMs are still 29.2 kOhm. The DC voltages out of the SRM satellite box are about 1.7 V. The signal was AC coupled using SR560 with two poles at 0.03 Hz and a gain of 10.

The noise is consistent with the one measured by the ADC except for the 3 Hz peak which does not show up in the ADC spectrum from Sunday. The peak appears in several channels I looked at. The instrument noise floor was measured by terminating the SR560 with 50 Ohm.

I recommend to change all OSEM transimpedance gains from 29 to 161 kV/A. Beyond this gain one will rail the AA filter module when the magnet is fully out of the OSEM.

The OSEM noise at 1 Hz is about factor of 10 above the shot noise. The damping loops impress this noise on the optics around the pendulum resonance frequency. Also the total contribution to the MC cavity length is sqrt(12) time the single sensor as there are 12 OSEMs contributing to MC length. The ADC noise is currently close but never the less not limiting the OSEM noise below 100 Hz. It can be further reduced by getting an extra factor of 2-3 in whitening gain above ~0.3 Hz. The rms of the ADC input of the modified PRM SD (R64 = 161 kOhm) channel is 10-20 cts during the day with damping loop off and whitening on.

The transimpedance amplifier LT1125CS is also not supposed to be limiting the noise. At 1 Hz the 1/f part of the noise: In<1pA/rtHz and Vn<20nV/rtHz.

Attachment 1: osemnoise.pdf  66 kB  | Hide | Hide all
osemnoise.pdf
ELOG V3.1.3-